at room temperature, an average hydrogen molecule travels at a speed of 1700 meters/second (about 3000 miles/hr) what is the speed of an average oxygen molecule under the same conditions?
1
Expert's answer
2020-12-08T05:33:29-0500

Average speed

"v = \sqrt{\frac{8RT}{πM}}"

R = gas constant

T = temperature

M = molar mass

Since R and T are constants:

"\frac{v}{v’} = \sqrt{\frac{M’}{M}}"

v = velocity of H2 = 1700 m/s

v’ = velocity of O2

M(H2) = 2 g/mol

M(O2) = 32 g/mol

"v’ = v \sqrt{\frac{M(H_2)}{M(O_2)}} \\

= 1700 \sqrt{\frac{2}{32}} \\

= 1700 \times 0.25 \\

= 425 \; m/s"

Answer: 425 m/s